# Principled RLHF from Heterogeneous Feedback via Personalization and Preference Aggregation

### v1

@@ 3 Personalized RLHF via Representation Learning
In this section, we provide the first approach in the personalization-based framework, based on representation learning.
@@
We will assume that we have access to a pre-trained feature function $\phi:{\mathcal{T}}\to\mathbb{R}^{d}$, which encodes a trajectory of states and actions (i.e., questions and answers) to a $d$-dimensional feature vector. This covers the case where feature $\phi_{h}:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}^{d}$ is defined at each state-action pair, i.e., $\phi(\tau):=\sum_{h\in[H]}\phi_{h}(s_{h},a_{h})$ for trajectory $\tau=(s_{1},a_{1},\dots,s_{H},a_{H})$. For example, it is common to use the penultimate layer of an existing pre-trained LLM or other pre-trained backbones to encode a long sentence to a feature vector (Donahue et al., 2014; Gulshan et al., 2016; Tang et al., 2015).
Our first goal is to learn reward models for each human user using preference datasets. First, we define the reward function class as
@@
3.1 Algorithms
We introduce our algorithms for learning personalized policy. Compared to traditional RLHF algorithms (Ziegler et al., 2019; Ouyang et al., 2022; Zhu et al., 2023), we consider personalized reward function by representation learning.
Algorithm 1 outputs a joint estimation of $\psi^{\star}$ and $\boldsymbol{\theta}^{\star}$ with maximum likelihood estimation (MLE), together with personalized policies. The input of the algorithm is $\widehat{\mathcal{D}}=\cup_{i\in[N]}\widehat{\mathcal{D}}_{i}$ where $\widehat{\mathcal{D}}_{i}=\{(o_{i}^{(j)},\tau_{i,0}^{(j)},\tau_{i,1}^{(j)})_{j\in[N_{p}]}\}$. Here, $\tau_{i,t}^{(j)}$ is sampled from the distribution $\mu_{t}$ for $t=0,1$, and $o_{i}^{(j)}\sim P_{r^{\star}_{i}}(\cdot|\tau_{0}^{(j)},\tau_{1}^{(j)})$. We regard the reward estimation of human users as a multi-task representation learning problem (Du et al., 2020; Tripuraneni et al., 2020). After estimating the reward functions, we construct two kinds of confidence sets for the reward function and outputs based on each confidence set:
  • Confidence set (Equation 3.1) for the MLE estimation as Liu et al. (2022), which is also used in (Liu et al., 2023; Zhan et al., 2023; Wang et al., 2024; Zhan et al., 2022), with $\zeta=C_{1}\log(\mathcal{N}_{\mathcal{G}_{\bm{r}}}(1/(NN_{p}))/\delta)$ for a constant $C_{1}>0$, which will be related to Theorem 3.1. the definition of bracketing number ($\mathcal{N}_{\mathcal{G}_{\bm{r}}}$) is deferred to Appendix B.
  • Confidence set (Equation 3.2) with $\zeta^{\prime}=C_{8}\left(k\frac{\xi^{2}\kappa^{2}\log(\mathcal{N}_{\mathcal{G}_{\bm{r}}}(1/(NN_{p}))/\delta)}{{\eta^{2}NN_{p}}}+\frac{\xi^{2}(k+\log(N/\delta))}{\eta^{2}N_{p}}+\lambda B^{2}\right)$, where $C_{8},\lambda>0$ are constants, $\xi:=\max_{x\in[-2R_{\max},2R_{\max}]}\left|\frac{\Phi^{\prime}(x)}{\Phi(x)}\right|$, $\kappa:=(\min_{x\in[-2R_{\text{max}},2R_{\text{max}}]}\Phi^{\prime}(x))^{-1}$, and $\eta:=\min_{x\in[-2R_{\max},2R_{\max}]}\left(\frac{\Phi^{\prime}(x)^{2}-\Phi^{\prime\prime}(x)\Phi(x)}{\Phi(x)^{2}}\right)$. In the case that $\Phi(x)=\sigma(x)$ (i.e. $\Phi$ is a Sigmoid function), $\xi\leq 1$ and $\kappa=\eta=\frac{1}{2+\exp(-2R_{\max})+\exp(2R_{\max})}$. This confidence set will be related to Theorem 3.4. With this confidence set, we can get sharper bound with Assumptions 3.2 - 3.4.
Lastly, we find the best policy based on the pessimistic expected value function. $\mu_{i,\text{ref}}$ in Algorithm 1 is a known reference trajectory distribution for individual $i\in[N]$, and it can be set as $\mu_{1}$.
@@ 3.1 Algorithms
$(\widehat{\omega},\widehat{\boldsymbol{\theta}})\leftarrow\operatorname*{arg\,min}_{\omega\in\Omega,\|\theta_{i}\|_{2}\leq B\text{ for all }i\in[N]}\sum_{i\in[N]}\sum_{j\in[N_{p}]}\log P_{\omega,\theta_{i}}(o_{i}^{(j)}\mid\tau_{i,0}^{(j)},\tau_{i,1}^{(j)})$
@@
$\widehat{\pi}_{i}\leftarrow\operatorname*{arg\,max}_{\pi\in\Pi}\min_{\bm{r}\in\mathcal{R}(\widehat{\mathcal{D}})}\left(J(\pi;r_{i})-\mathbb{E}_{\tau\sim\mu_{i,\text{ref}}}[r_{i}(\tau)]\right)$ (3.3)
or
@@
  Output: $(\widehat{\omega},\widehat{\boldsymbol{\theta}},(\widehat{\pi}_{i})_{i\in[N]})$ or $(\widehat{\omega},\widehat{\boldsymbol{\theta}},(\widehat{\pi}^{\prime}_{i})_{i\in[N]})$.
@@
$\widehat{\theta}_{0}\leftarrow\operatorname*{arg\,min}_{\|\theta_{0}\|_{2}\leq B}\sum_{j\in[N_{p}]}\log P_{\widehat{\omega},\theta_{0}}(o_{0}^{(j)}\mid\tau_{0,0}^{(j)},\tau_{0,1}^{(j)})$
@@
3.2 Results and Analyses
For ease of analysis, we consider the case where the sizes of preference datasets for each individual $i\in\{0\}\cup[N]$ are identical, i.e., $\widehat{\mathcal{D}}_{i}=\{(o_{i}^{(j)},\tau_{i,0}^{(j)},\tau_{i,1}^{(j)})_{j\in[N_{p}]}\}$, satisfies $|\widehat{\mathcal{D}}_{i}|=N_{p}$ for all $i\in\{0\}\cup[N]$. The result in this section can also be extended to the case with $|\widehat{\mathcal{D}}_{i}|=N_{p,i}$ for each individual $i$. We defer all the proofs of this section to Appendix C.
@@ 3.2 Results and Analyses
We also define the concentrability coefficient of the reward scalar class in Section B.2, and we denote this as $C_{r}(\mathcal{G}_{r},\pi_{\text{tar}},\mu_{\text{ref}})$.
@@
We present the gap of the expected value function between the target policy $\pi_{i,\text{tar}}$ and the estimated policy $\widehat{\pi}_{i}$ for each individual $i\in[N]$. Here, $\pi_{i,\text{tar}}$, which may be the optimal policy $\pi_{i}^{\star}$ over $r_{i}^{\star}$, serves as the policy that $\widehat{\pi}_{i}$ will compete.
Theorem 3.1.
(Total Expected Value Function Gap). Suppose 3.1 holds. For any $\delta\in(0,1]$, with probability at least $1-\delta$, the output $(\widehat{\pi}_{i})_{i\in[N]}$ of Algorithm 1 satisfies
Corollary 3.2.
(Expected Value Function Gap). Suppose 3.1 holds. For any $\delta\in(0,1]$ and all $i\in[N]$, with probability at least $1-\delta$, the output $\widehat{\pi}_{i}$ of Algorithm 1 satisfies
where $c>0$ is a constant.
Note that the results above do not need any assumption on $(\theta_{i}^{\star})_{i\in[N]}$. Still, as $N_{p}\to\infty$, $\widehat{\pi}_{i}$ has comparable or better performance than the comparator policy $\pi_{i,\text{tar}}$, which approaches the optimal policy if $\pi_{i,\text{tar}}=\pi_{i}^{\star}$. We defer the proofs of Theorem 3.1 and Corollary 3.2 to Section C.1.
Next, we want to improve the bound for Corollary 3.2, as the gap of the expected value function does not decay with $N$, which is the number of human users. Therefore, we consider the case that $(\theta_{i})_{i\in[N]}$ are diverse (3.2), which is critical for improving the sample complexity of Algorithm 1 by outputting $(\widehat{\pi}_{i}^{\prime})_{i\in[N]}$. We will additionally assume the uniqueness of the representation up to the orthonormal linear transformation (3.3), and uniform concentration of covariance (3.4). These assumptions are commonly used in multi-task learning (Du et al., 2020; Tripuraneni et al., 2021; Lu et al., 2021)
Assumption 3.2 (Diversity).
@@
3.2 means that $\theta_{i}$ is evenly distributed in $\mathbb{R}^{d}$ space for $i\in[N]$, which indicates “diverse” human reward function.
Assumption 3.3 (Uniqueness of Representation (up to Orthonormal-Transformation)).
For any representation functions $\psi,\psi^{\prime}\in\Psi$ and $\epsilon>0$, if there exists $\{v_{i}\}_{i=1}^{T},\{v_{i}^{\prime}\}_{i=1}^{T}$, and a trajectory distribution $\mu$ that satisfy
$\frac{1}{T}\sum_{i\in[T]}\mathbb{E}_{\tau\sim\mu}\|\psi(\phi(\tau))^{\top}v_{i}-\psi^{\prime}(\phi(\tau))^{\top}v_{i}^{\prime}\|^{2}\leq\epsilon,$
$W=[v_{1},v_{2},\cdots,v_{T}]\in\mathbb{R}^{k\times T}$ satisfies $\sigma^{2}_{k}(W)\geq\Omega\left(T/k\right)$, and $\|v_{i}\|_{2}\leq B$ for all $i\in[T]$. Then, there exists a constant orthonormal matrix $P$ such that
@@
This assumption posits that if two representation functions, $\psi$ and $\psi^{\prime}$, yield sufficiently small differences in expected squared norms of their inner products with corresponding vectors over trajectory distributions, then they are related by a constant orthonormal transformation. If $\psi_{\omega}(\phi(s,a)):=\omega\phi(s,a)$ where $\omega$ is ${k\times d}$ orthonormal matrix, we can prove that 3.3 holds with non-degenerate $\phi(s,a)$ distribution (Section C.2.2).
@@
Assumption 3.4.
(Uniform Concentrability). For any $\delta\in(0,1]$, there exists a number $N_{\text{unif}}(\Psi,\mu_{0},\mu_{1},\delta)$ such that for any $n\geq N_{\text{unif}}(\Psi,\mu_{0},\mu_{1},\delta)$, the empirical estimation $\widehat{\Lambda}_{\psi,\psi^{\prime}}(\mu_{0},\mu_{1})$ of $\Lambda_{\psi,\psi^{\prime}}(\mu_{0},\mu_{1})$ based on $n$ independent trajectory sample pairs from distributions $(\mu_{0},\mu_{1})$, with probability at least $1-\delta$, will satisfy the following inequality:
$1.1\Lambda_{\psi,\psi^{\prime}}(\mu_{0},\mu_{1})\succeq\widehat{\Lambda}_{\psi,\psi^{\prime}}(\mu_{0},\mu_{1})\succeq 0.9\Lambda_{\psi,\psi^{\prime}}(\mu_{0},\mu_{1}),$
for all $\psi,\psi^{\prime}\in\Psi$.
3.4 means that the empirical estimate $\widehat{\Lambda}_{\psi,\psi^{\prime}}(\mu_{0},\mu_{1})$ closely approximates the true $\Lambda_{\psi,\psi^{\prime}}(\mu_{0},\mu_{1})$ with high probability. Similarly, if $\psi_{\omega}(\phi(\tau)):=\omega\phi(\tau)$, $N_{\text{point}}(\Psi,\mu_{0},\mu_{1},\delta)=\widetilde{\mathcal{O}}(d)$ (Du et al., 2020, Claim A.1). If distributions $\mu_{0},\mu_{1}$ are clear from the context, we omit the notation $\mu_{0},\mu_{1}$ for $\Sigma_{\psi,\psi^{\prime}}(\mu_{0},\mu_{1})$ and $\Lambda_{\psi,\psi^{\prime}}(\mu_{0},\mu_{1})$. Moreover, we also write $\Sigma_{\psi,\psi}$ as $\Sigma_{\psi}$ for notational convenience.
With 3.2 and 3.3, $\psi^{\star}$ and $\psi_{\omega}$ are close up to an orthonormal matrix transformation, as asserted below:
@@
(Closeness between $\psi^{\star}$ and $\psi_{\omega}$). Suppose Assumptions 3.1, 3.2, and 3.3 hold. For any $\delta\in(0,1]$, with probability at least $1-\delta$, if $\bm{r}_{\omega,\boldsymbol{\theta}}\in\mathcal{R}^{\prime}(\mathcal{D})$ as specified in Algorithm 1, then there exists an orthonormal matrix $P_{\omega}$ such that
@@
We provide the proof of Corollary 3.3 in Section C.3. Using Corollary 3.3, now we have a better sample complexity compared with that in Theorem 3.1, as stated below:
@@
(Improved Expected Value Function Gap). Suppose Assumptions 3.1, 3.2, 3.3, and 3.4 hold. For any $\delta\in(0,1]$, all $i\in[N]$ and $\lambda>0$, with probability at least $1-\delta$, the output $\widehat{\pi}_{i}^{\prime}$ of Algorithm 1 satisfies
@@
(Expected Value Function Gap for a New Human User). Suppose Assumptions 3.2, 3.3, and 3.4 hold. For any $\delta\in(0,1]$ and $\lambda>0$, with probability at least $1-\delta$, the output $\widehat{\pi}_{0}$ of Algorithm 2 satisfies
@@
provides $\mathcal{N}_{\mathcal{G}_{\bm{r}}^{\prime}}(1/(NN_{p}))\leq\mathcal{O}\left(Nd\log\left(R_{\max}NN_{p}/\delta\right)\right)$. Since $d\gg k$, the bound of Equation 3.5’s right-hand side has a significant improvement when we use the representation learning. If the representation function class is an MLP class, we can use a known bracket number by Bartlett et al. (2017).
We also point out that the existing technique from representation learning literature does not cover the case with general representation function learning with a log-likelihood loss function with $\mathcal{O}(1/N_{p})$ rate, to the best of our knowledge.
@@
${\rm CB}(\Lambda)\coloneqq\left\{Q\coloneqq\left(\left\{\mu_{0},\mu_{1}\right\},\{\tau_{i,0}^{(j)},\tau_{i,1}^{(j)}\}_{i\in[N],j\in[N_{p}]},\omega,\boldsymbol{\theta}\right){\,\big{|}\,}C_{\bm{r}}(\mathcal{G}_{\bm{r}},\pi^{\star},\mu_{1},i)\leq\Lambda\text{ for all }i\in[N]\right\}$
is the family of MDP with $N$ reward functions and $H=1$ instances.
The proof is largely adapted from (Zhu et al., 2023, Theorem 3.10). By Theorem 3.6, for general representation function class, we establish that Algorithm 1 is near-optimal for the sub-optimality of the induced personalization policy, as $\log(\mathcal{N}_{\mathcal{G}_{\bm{r}}}(1/NN_{p}))$ can be small so that $\sqrt{k\frac{\log(\mathcal{N}_{\mathcal{G}_{\bm{r}}}(1/(NN_{p}))/\delta)}{NN_{p}}}$ can be dominated by $\sqrt{\frac{k}{N_{p}}}$. Note that if $\Psi$ is a linear representation class, our result for personalization (Theorem 3.4) still has a $\sqrt{k}$ gap compared to the lower bound (Theorem 3.6). This gap is also observed in (Tripuraneni et al., 2020). We will leave the sharpening of this $\sqrt{k}$ factor as a future work.
4 Personalized RLHF via Human User Clustering
We now provide the second approach in the personalization-based framework, through human user clustering. In particular, fine-tuning an LLM for each individual may be impractical. We thus propose an alternative approach that segments human users into clusters and fine-tunes an LLM for each cluster. This strategy entails deploying $K$ clustered models, which can be smaller than the number of human users $N$. A critical aspect of this methodology is the way to generate clusters. This clustering-based personalization has also been studied in the federated (supervised) learning literature (Mansour et al., 2020; Ghosh et al., 2020; Sattler et al., 2020). We introduce our algorithm next, based on the algorithmic idea in Mansour et al. (2020).
4.1 Algorithms

### v2

@@ 3.1 Learning Personalized Reward Model
In this subsection, we provide the first approach in the personalization-based framework, based on representation learning.
@@
We will assume that we have access to a pre-trained feature function $\phi:{\mathcal{T}}\to\mathbb{R}^{d}$, which encodes a trajectory of states and actions (i.e., questions and answers) to a $d$-dimensional feature vector. This covers the case where feature $\phi_{h}:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}^{d}$ is defined at each state-action pair, i.e., $\phi(\tau):=\sum_{h\in[H]}\phi_{h}(s_{h},a_{h})$ for trajectory $\tau=(s_{1},a_{1},\dots,s_{H},a_{H})$. For example, it is common to use the penultimate layer of an existing pre-trained LLM or other pre-trained backbones to encode a long sentence to a feature vector (Donahue et al., 2014; Gulshan et al., 2016; Tang et al., 2016).
Our first goal is to learn multiple reward models for each human user using preference datasets. First, we define the reward function class as
@@
3.1.1 Algorithms
We introduce our algorithm for learning personalized policy. Compared to traditional RLHF algorithms (Ziegler et al., 2019; Ouyang et al., 2022; Zhu et al., 2023), we consider personalized reward function by representation learning.
Algorithm 1 outputs a joint estimation of $\psi^{\star}$ and $\boldsymbol{\theta}^{\star}$ with maximum likelihood estimation (MLE), together with personalized policies. The input of the algorithm is $\widehat{\mathcal{D}}=\cup_{i\in[N]}\widehat{\mathcal{D}}_{i}$ where $\widehat{\mathcal{D}}_{i}=\{(o_{i}^{(j)},\tau_{i,0}^{(j)},\tau_{i,1}^{(j)})_{j\in[N_{p}]}\}$. Here, $\tau_{i,t}^{(j)}$ is sampled from the distribution $\mu_{t}$ for $t=0,1$, and $o_{i}^{(j)}\sim P_{r^{\star}_{i}}(\cdot|\tau_{0}^{(j)},\tau_{1}^{(j)})$. First, we estimate the reward function of human users. After estimating the reward functions, we construct a confidence set for the reward function as follows: Confidence set (Equation 3.1) with $\zeta^{\prime}=C_{8}\left(k\frac{\xi^{2}\kappa^{2}\log(\mathcal{N}_{\mathcal{G}_{\bm{r}}}(1/(NN_{p}))/\delta)}{{\eta^{2}NN_{p}}}+\frac{\xi^{2}(k+\log(N/\delta))}{\eta^{2}N_{p}}+\lambda B^{2}\right)$, where $C_{8},\lambda>0$ are constants, $\xi:=\max_{x\in[-2R_{\max},2R_{\max}]}\left|\frac{\Phi^{\prime}(x)}{\Phi(x)}\right|$, $\kappa:=(\min_{x\in[-2R_{\text{max}},2R_{\text{max}}]}\Phi^{\prime}(x))^{-1}$, and $\eta:=\min_{x\in[-2R_{\max},2R_{\max}]}\left(\frac{\Phi^{\prime}(x)^{2}-\Phi^{\prime\prime}(x)\Phi(x)}{\Phi(x)^{2}}\right)$. In the case that $\Phi(x)=\sigma(x)$ (i.e. $\Phi$ is a Sigmoid), $\xi\leq 1$ and $\kappa=\eta=\frac{1}{2+\exp(-2R_{\max})+\exp(2R_{\max})}$. This confidence set will be related to Theorem 3.1. Lastly, we find the best policy based on the pessimistic expected value function. $\mu_{i,\text{ref}}$ in Algorithm 1 is a known reference trajectory distribution for individual $i\in[N]$, and it can be set as $\mu_{1}$.
@@ 3.1.1 Algorithms
$(\widehat{\omega},\widehat{\boldsymbol{\theta}})\leftarrow\operatorname*{arg\,max}_{\omega\in\Omega,\|\theta_{i}\|_{2}\leq B\text{ for all }i\in[N]}\sum_{i\in[N]}\sum_{j\in[N_{p}]}\log P_{\omega,\theta_{i}}(o_{i}^{(j)}\mid\tau_{i,0}^{(j)},\tau_{i,1}^{(j)})$
@@
  Output: $(\widehat{\omega},\widehat{\boldsymbol{\theta}},(\widehat{\pi}^{\prime}_{i})_{i\in[N]})$.
@@
$\widehat{\theta}_{0}\leftarrow\operatorname*{arg\,max}_{\|\theta_{0}\|_{2}\leq B}\sum_{j\in[N_{p}]}\log P_{\widehat{\omega},\theta_{0}}(o_{0}^{(j)}\mid\tau_{0,0}^{(j)},\tau_{0,1}^{(j)})$
@@
3.1.2 Results and Analyses
For ease of analysis, we consider the case where the sizes of preference datasets for each individual $i\in\{0\}\cup[N]$ are identical, i.e., $\widehat{\mathcal{D}}_{i}=\{(o_{i}^{(j)},\tau_{i,0}^{(j)},\tau_{i,1}^{(j)})_{j\in[N_{p}]}\}$, satisfies $|\widehat{\mathcal{D}}_{i}|=N_{p}$ for all $i\in\{0\}\cup[N]$. The result in this section can also be extended to the case with $|\widehat{\mathcal{D}}_{i}|=N_{p,i}$ for each individual $i$. We defer all the proofs of this section to Appendix F.
@@ 3.1.2 Results and Analyses
We also define the concentrability coefficient of the reward scalar class in Appendix D, and we denote this as $C_{r}(\mathcal{G}_{r},\pi_{\text{tar}},\mu_{\text{ref}})$.
@@
We consider the case that $(\theta_{i})_{i\in[N]}$ are diverse (2), which is critical for improving the sample complexity of Algorithm 1 by outputting $(\widehat{\pi}_{i}^{\prime})_{i\in[N]}$. We will additionally assume the uniqueness of the representation up to the orthonormal linear transformation (3), and uniform concentration of covariance (4). These assumptions are commonly used in multi-task learning (Du et al., 2021; Tripuraneni et al., 2021; Lu et al., 2021)
Assumption 2 (Diversity).
@@
2 means that $\theta_{i}$ is evenly distributed in $\mathbb{R}^{d}$ space for $i\in[N]$, which indicates “diverse” human reward function.
Assumption 3 (Uniqueness of Representation (up to Orthonormal-Transformation)).
For any representation functions $\psi,\psi^{\prime}\in\Psi$ and $\epsilon>0$, if there exists $\{v_{i}\}_{i=1}^{T},\{v_{i}^{\prime}\}_{i=1}^{T}$, and a trajectory distribution $\mu$ that satisfy $\frac{1}{T}\sum_{i\in[T]}\mathbb{E}_{\tau\sim\mu}\|\psi(\phi(\tau))^{\top}v_{i}-\psi^{\prime}(\phi(\tau))^{\top}v_{i}^{\prime}\|^{2}\leq\epsilon,$ where $W=[v_{1},v_{2},\cdots,v_{T}]\in\mathbb{R}^{k\times T}$ satisfies $\sigma^{2}_{k}(W)\geq\Omega\left(T/k\right)$, and $\|v_{i}\|_{2}\leq B$ for all $i\in[T]$. Then, there exists a constant orthonormal matrix $P$ such that
@@
This assumption posits that if two representation functions, $\psi$ and $\psi^{\prime}$, yield sufficiently small differences in expected squared norms of their inner products with corresponding vectors over trajectory distributions, then they are related by a constant orthonormal transformation. If $\psi_{\omega}(\phi(s,a)):=\omega\phi(s,a)$ where $\omega$ is ${k\times d}$ orthonormal matrix, we can prove that 3 holds with non-degenerate $\phi(s,a)$ distribution (Section F.3.2).
@@
Assumption 4.
(Uniform Concentrability). For any $\delta\in(0,1]$, there exists a number $N_{\text{unif}}(\Psi,\mu_{0},\mu_{1},\delta)$ such that for any $n\geq N_{\text{unif}}(\Psi,\mu_{0},\mu_{1},\delta)$, the empirical estimation $\widehat{\Lambda}_{\psi,\psi^{\prime}}(\mu_{0},\mu_{1})$ of $\Lambda_{\psi,\psi^{\prime}}(\mu_{0},\mu_{1})$ based on $n$ independent trajectory sample pairs from distributions $(\mu_{0},\mu_{1})$, with probability at least $1-\delta$, will satisfy the following inequality for all $\psi,\psi^{\prime}\in\Psi$:
$1.1\Lambda_{\psi,\psi^{\prime}}(\mu_{0},\mu_{1})\succeq\widehat{\Lambda}_{\psi,\psi^{\prime}}(\mu_{0},\mu_{1})\succeq 0.9\Lambda_{\psi,\psi^{\prime}}(\mu_{0},\mu_{1}).$
4 means that the empirical estimate $\widehat{\Lambda}_{\psi,\psi^{\prime}}(\mu_{0},\mu_{1})$ closely approximates the true $\Lambda_{\psi,\psi^{\prime}}(\mu_{0},\mu_{1})$ with high probability. Similarly, if $\psi_{\omega}(\phi(\tau)):=\omega\phi(\tau)$, $N_{\text{point}}(\Psi,\mu_{0},\mu_{1},\delta)=\widetilde{\mathcal{O}}(d)$ (Du et al., 2021, Claim A.1). If distributions $\mu_{0},\mu_{1}$ are clear from the context, we omit the notation $\mu_{0},\mu_{1}$ for $\Sigma_{\psi,\psi^{\prime}}(\mu_{0},\mu_{1})$ and $\Lambda_{\psi,\psi^{\prime}}(\mu_{0},\mu_{1})$. Moreover, we also write $\Sigma_{\psi,\psi}$ as $\Sigma_{\psi}$ for notational convenience.
With 2 and 3, $\psi^{\star}$ and $\psi_{\omega}$ are close up to an orthonormal matrix transformation, as asserted below:
@@
(Closeness between $\psi^{\star}$ and $\psi_{\omega}$). Suppose Assumptions 1, 2, and 3 hold. For any $\delta\in(0,1]$, with probability at least $1-\delta$, if $\bm{r}_{\omega,\boldsymbol{\theta}}\in\mathcal{R}^{\prime}(\mathcal{D})$ as specified in Algorithm 1, then there exists an orthonormal matrix $P_{\omega}$ such that
@@
We present the gap of the expected value function between the target policy $\pi_{i,\text{tar}}$ and the estimated policy $\widehat{\pi}_{i}$ for each individual $i\in[N]$. Here, $\pi_{i,\text{tar}}$, which may be the optimal policy $\pi_{i}^{\star}$ over $r_{i}^{\star}$, serves as the policy that $\widehat{\pi}_{i}$ will compare with.
@@
(Expected Value Function Gap). Suppose Assumptions 1, 2, 3, and 4 hold. For any $\delta\in(0,1]$, all $i\in[N]$ and $\lambda>0$, with probability at least $1-\delta$, the output $\widehat{\pi}_{i}^{\prime}$ of Algorithm 1 satisfies
@@
(Expected Value Function Gap for a New Human User). Suppose Assumptions 2, 3, and 4 hold. For any $\delta\in(0,1]$ and $\lambda>0$, with probability at least $1-\delta$, the output $\widehat{\pi}_{0}$ of Algorithm 2 satisfies
@@
provides $\mathcal{N}_{\mathcal{G}_{\bm{r}}^{\prime}}(1/(NN_{p}))\leq\mathcal{O}\left(Nd\log\left(R_{\max}NN_{p}/\delta\right)\right)$. Since $d\gg k$, the bound of Equation 3.3’s right-hand side has a significant improvement when we use representation learning. If the representation function class is an MLP class, we can use a known bracket number by (Bartlett et al., 2017).
We also point out that the existing technique from representation learning literature does not cover the case with general representation function learning with a log-likelihood loss function with $\mathcal{O}(1/N_{p})$ rate, to the best of our knowledge. The technical results are thus of independent interest.
@@
${\rm CB}(\Lambda)\coloneqq\left\{Q\coloneqq\left(\left\{\mu_{0},\mu_{1}\right\},\{\tau_{i,0}^{(j)},\tau_{i,1}^{(j)}\}_{i\in[N],j\in[N_{p}]},\omega,\boldsymbol{\theta}\right){\,\big{|}\,}C_{\bm{r}}^{\prime}(\mathcal{G}_{\bm{r}},\pi^{\star},\mu_{1},i)\leq\Lambda\text{ for all }i\in[N]\right\}$
is the family of MDP with $N$ reward functions and $H=1$ instances, where
$C_{\bm{r}}^{\prime}(\mathcal{G}_{\bm{r}},\pi^{\star},\mu_{1},i)\coloneqq\max\left\{0,\sup_{\bm{r}\in\mathcal{G}_{\bm{r}}}\frac{\mathbb{E}_{\tau_{0}\sim\pi^{\star},\tau_{1}\sim\mu_{1}}\left[r^{\star}_{i}\left(\tau_{0}\right)-r^{\star}_{i}\left(\tau_{1}\right)-r_{i}\left(\tau_{0}\right)+r_{i}\left(\tau_{1}\right)\right]}{\sqrt{\frac{1}{N_{p}}\sum_{j=1}^{N_{p}}\left[\left|r^{\star}_{i}\left(\tau_{i,0}^{(j)}\right)-r^{\star}_{i}\left(\tau_{i,1}^{(j)}\right)-r_{i}\left(\tau_{i,0}^{(j)}\right)+r_{i}\left(\tau_{i,1}^{(j)}\right)\right|^{2}\right]}}\right\}.$ (3.4)
Our approach for personalized reward lower bound builds upon (Zhu et al., 2023, Theorem 3.10). Note that all results in this paper still hold for the new concentrability coefficient $C_{\bm{r}}^{\prime}$. By Theorem 3.3, for general representation function class, we establish that Algorithm 1 is near-optimal for the sub-optimality of the induced personalization policy, as $\log(\mathcal{N}_{\mathcal{G}_{\bm{r}}}(1/NN_{p}))$ can be small so that $\sqrt{k\frac{\log(\mathcal{N}_{\mathcal{G}_{\bm{r}}}(1/(NN_{p}))/\delta)}{NN_{p}}}$ can be dominated by $\sqrt{\frac{k}{N_{p}}}$. Note that if $\Psi$ is a linear representation class, our result for personalization (Theorem 3.1) still has a $\sqrt{k}$ gap compared to the lower bound (Theorem 3.3). This gap is also observed in (Tripuraneni et al., 2020). We will leave the sharpening of this $\sqrt{k}$ factor as a future work.
3.2 Personalized RLHF via Human User Clustering
We now provide the second approach in the personalization-based framework, through human user clustering. In particular, fine-tuning an LLM for each individual may be impractical. We thus propose an alternative approach that segments human users into clusters and fine-tunes an LLM for each cluster. This strategy entails deploying $K$ clustered models, which can be smaller than the number of human users $N$. A critical aspect of this methodology is the way to generate clusters. This clustering-based personalization has also been studied in the federated (supervised) learning literature (Mansour et al., 2020; Ghosh et al., 2020; Sattler et al., 2020). We introduce our algorithm next, based on the algorithmic idea in (Mansour et al., 2020).
3.2.1 Algorithms